A Borg-type theorem for Super Symmetric Dirac Difference Operators

Stephen Clark University of Missouri-Rolla, USA

We discuss a Borg-type uniqueness theorem for matrix-valued super symmetric Dirac difference operators. This is joint work with Fritz Gesztesy of the University of Missouri-Columbia.